find three consecutive integers such that the product of the first and second is five less than five times the third
Answer by richwmiller(17219)

n*(n+1)=5*(n+2)-5
n^2+n=5n+10-5
n^2-4n-5=0
n=-1 n=5
check
-1*0=5(1)-5
0=0
5*6=5*7-5
30=30
ok
-1,0,1
and
5,6,7